[Hallc_running] [New Logentry] Status Update

pooser at jlab.org pooser at jlab.org
Sun Feb 25 13:50:01 EST 2018

Logentry Text:
Beam during swing yesterday and owl this morning was relatively stable at 40-45 uA while maintaining an average trip rate of ~10 trips/hr (https://logbooks.jlab.org/entry/3538313) resulting in a sizable amount of production data collected.   We are still on target to finish the SHMS 39 degrees running at -2.5 GeV by day shift tomorrow and maybe even as soon as start of owl tonight, fingers crossed!  We have currently obtained ~78% of the LH2 data, ~71% of the LD2 data, ~72% of the 10 cm Al dummy data (https://docs.google.com/document/d/13i-SQGGES52d8fjJo9CTdgMc0A7J7z_t9FdKewW6HEM/edit).   The next kinematic point for the SHMS is to remain at 39 degrees while the momentum is lowered to -2.0 Gev (https://hallcweb.jlab.org/DocDB/0009/000953/001/F2_EMC_runlist_final.pdf).

During the beginning of owl, the shift crew encountered TDC1190 errors during the replay of HMS run 1680 (https://logbooks.jlab.org/entry/3538054).  This error is indicative of the HMS chambers ringing.  The shift crew had to execute “kcoda” then “startcoda” twice before the issue was resolved.  The usual prescription is to raise the DC threshold to -20 V for about a minute or so, and then lower the threshold back to the nominal value of -5.5 V (https://logbooks.jlab.org/entry/3534659).  In any case, the issue was quickly resolved and minimal time was lost.

This morning the liquid Helium levels in HMS Q1 began to drop (https://logbooks.jlab.org/entry/3538237) as a result of the top fill valve (JT8) closing.  The suspicion is that the 24 V relay chip failed in a closed position and forced the valve to close and settle to its mechanical lower limit (https://logbooks.jlab.org/entry/3538239).  Thanks to the quick response of Simona and Steve L., the LHe fill flow was switched to the bottom fill valve (JT6) and the LHe level began to rise to nominal levels.  During this time the JT8 valve came back to life and is now currently maintaining the LHe levels.  Shift crew should remain diligent and keep a close eye on Q1’s LHe levels.

MCC will begin beam studies tomorrow morning at 9 am however, we will take beam parasitically until 1 pm.  With MCC planning to conclude beam studies at 4 pm, any work that is planned in the hall must take less than 3 hours.  If you plan to conduct any work in the all tomorrow please verify with me (pooser at jlab.org) that you will be able to start this work promptly at 1 pm and that it will conclude no later than 4 pm.

Work currently planned during Monday’s (2/26) beam studies starting at 1 pm:
Joe Beaufait: diagnose (and potentially fix) HMS door/shutter breaker (https://logbooks.jlab.org/entry/3535292)
Andrew Kenyon: install five Y strainers on the SHMS power supply LCW lines.  The power supplies will have to be ramped down and turned off
Howard Fenker: visually inspect HMS hut for causes for chamber ringing and 60 Hz noise

Special notices in effect for shift crews:
Trouble shooting HMS Q1 valves JT6 and JT8: https://logbooks.jlab.org/entry/3538239, https://logbooks.jlab.org/entry/3538243, https://logbooks.jlab.org/entry/3538254
Trouble shooting HMS Q2 power supply: https://logbooks.jlab.org/entry/3537306
Don't ramp HMS Q2 higher than 610A: https://logbooks.jlab.org/entry/3535714
TO Instructions for LH2 <--> LD2 switches : https://logbooks.jlab.org/entry/3534487


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3538355
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180225/2ff4830e/attachment.html>

More information about the Hallc_running mailing list